McDougan Associates​ (U.S.). McDougan​ Associates, a​ U.S.-based investment​ partnership, borrows €85,000,000 at a time when the exchange rate is $1.3395​/€. The entire principal is to be repaid in three​ years, and interest is 6.250​% per​ annum, paid annually in euros. The euro is expected to depreciate​ vis-à-vis the dollar at 3.3​% per annum. What is the effective cost of this loan for​ McDougan?

Complete the following table to calculate the dollar cost of the​ euro-denominated debt for years 0 through 3. Enter a positive number for a cash inflow and negative for a cash outflow.  ​(Round the amount to the nearest whole number and the exchange rate to four decimal​ places.)

Year 0 Year 1 Year 2 Year 3

Proceeds from borrowing euros

€ 85,000,000

Interest payment due in euros

Repayment of principal in year 3 -85,000,000
Total cash flow of euro-dominated debt

Expected exchange rate, $/€

1.3395

Dollar equivalent of euro-denominated

$ $ $ $

                           cash flow
Answer #1

Answer:

Sample calculation:

Interest payment for each year= 6.25%*85000000=5,312,500

dollar Equivalent of euro denominated for year 0=$1.3395

Depreciation rate =3.3%

dollar Equivalent of euro denominated for year 1=$1.3395/(1-3.3%)=$1.3852

dollar Equivalent of euro denominated for year 1=$1.3852/(1-3.3%)=$1.4325

dollar Equivalent of euro denominated for year 1=$1.4325/(1-3.3%)=$1.4814

Year 0 Year 1 Year 2 Year 3
Proceeds from Borrowing funds 85,000,000
Interest Rate due in euros -5312500 -5312500 -5312500
Repayment of principal in year 3 -85000000
Total cash flow of euro dominated debt in £ 85000000 -5312500 -5312500 -90312500
Expected Exchange rate, $/£ 1.3395
Dollar equivalent of euro denominated in $ $1.3395 $1.3852 $1.4325 $1.4814
Cash flow $113857500 -$7358938.728 -$7610071.073 -$133786151.2